OpenStudy (anonymous):

What is the equation in standard form of the line which passes through (-2, -3) and has a slope of 1/2? Give a medal to first that answers

OpenStudy (anonymous):

@Mertsj

OpenStudy (mertsj):

First write the equation of the line in point slope form and then rearrange it.

OpenStudy (anonymous):

ok one sec

OpenStudy (mertsj):

Point slope form is: \[y-y _{1}=m(x-x _{1})\]

OpenStudy (anonymous):

y-(-3)=1/2x-(-1

OpenStudy (anonymous):

i have a few others i need help with

OpenStudy (anonymous):

i did that what is next i put it in to standard form but it never matches up with the answers they give you

OpenStudy (mertsj):

\[y+2=\frac{1}{2}(x+2)\]

OpenStudy (anonymous):

how did you get 2, the y value is -3

OpenStudy (anonymous):

can you just do all the steps up to the last one and i can find the answer then

OpenStudy (mertsj):

typo

OpenStudy (anonymous):

ok thats what i thought

OpenStudy (mertsj):

\[y+3=\frac{1}{2}(x+2)\] \[y+3=\frac{1}{2}x +1\] \[\frac{1}{2}x-y=2\]

OpenStudy (mertsj):

Or sometimes they want it without fractions in which case it would be: \[x-2y=4\]

OpenStudy (anonymous):

so the answer is x-2y=4

OpenStudy (anonymous):

To which graph does the point (2, -3) belong? Choose one answer. a. y > -x - 2 b. y > -x - 1 c. y > -2x + 3 d. y > -3x + 5 here is another i need help with

OpenStudy (mertsj):

Did you plug in 2 for x and -3 for y and see which one is true?

OpenStudy (anonymous):

i don't know what i am looking for though

OpenStudy (mertsj):

-3>-2-2, -3>-4 Well shazam!! the very first one is true.

OpenStudy (anonymous):

ha ha i was right that is what i thought it was 2 more

OpenStudy (anonymous):

What is the equation of the line in standard form that passes through the point (3, -2) and is parallel to the line y = -4x + 1?

OpenStudy (mertsj):

Do you know that parallel lines have the same slope?

OpenStudy (mertsj):

So this is just like the first one with slope = -4

OpenStudy (anonymous):

so it would look like y-(-2)=-4(x-3)

OpenStudy (mertsj):

\[y+2=-4(x-3)\]

OpenStudy (mertsj):

yep. Now rearrange it.

OpenStudy (anonymous):

-5x+y=14 is that the answer

OpenStudy (mertsj):

\[y+2=-4x+12\] \[4x+y=10\]

OpenStudy (anonymous):

one more

OpenStudy (anonymous):

acctually 2 more

OpenStudy (anonymous):

What is the equation of the line in slope-intercept form that passes through the point (3, -1) and is perpendicular to the line y = x + 1?

OpenStudy (mertsj):

What is the slope of y = x+1?

OpenStudy (anonymous):

1

OpenStudy (anonymous):

\[y+1=2x-3\]

OpenStudy (anonymous):

\[y=2x-4\]

OpenStudy (anonymous):

is that the answer

OpenStudy (mertsj):

You didn't answer my question. what is the slope of y=x+1?

OpenStudy (anonymous):

1

OpenStudy (mertsj):

Then what is the slope of the perpendicular?

OpenStudy (mertsj):

Hint. It is the negative reciprocal of 1

OpenStudy (anonymous):

-1

OpenStudy (mertsj):

ok. Now use point slope form and plug in the given point and the slope -1

OpenStudy (anonymous):

alright one sec

OpenStudy (anonymous):

y+1=-2x+3

OpenStudy (mertsj):

\[y+1=-1(x-3)\] y+1=-x+3 \[y=-x+2\]

OpenStudy (anonymous):

y=x+2 is the answer

OpenStudy (mertsj):

Did you notice the negative sign in front of the x?

OpenStudy (anonymous):

What is the vertex of the graph of y = |x - 2|- 5? last one

OpenStudy (anonymous):

yes typo

OpenStudy (mertsj):

Remember...the slope of the perpendicular is -1

OpenStudy (mertsj):

(2,-5)

OpenStudy (anonymous):

one more

OpenStudy (anonymous):

never mind thank you so much